OpenStudy (anonymous):

Integral of: (5x-1)/(2x^2-x)dx I'm trying to use PF but something is not right. :c

OpenStudy (anonymous):

I get the answer: lnx +3ln(2x-1)

OpenStudy (anonymous):

Where the answer should be: lnx +3/2ln(2x-1) according to book n' wolf.

OpenStudy (turingtest):

\[{5x-1\over2x^2-x}=\frac Ax+\frac B{2x-1}\]is this how you started?

OpenStudy (anonymous):

Indeed!

OpenStudy (anonymous):

first decompose it to \[\int\limits_{}^{}1/x+3/2x-1dx\]

OpenStudy (turingtest):

well then you probable made some small error\[{5x-1\over2x^2-x}=\frac Ax+\frac B{2x-1}\implies A(2x-1)+Bx=5x-1\]choose x=1/2 and x=0 and see what you get

OpenStudy (anonymous):

then integrate to end up with \[\ln \left| x \right|+(3/2)\ln \left| 2x-1 \right|+c\]

OpenStudy (turingtest):

\[{5x-1\over2x^2-x}=\frac Ax+\frac B{2x-1}\implies A(2x-1)+Bx=5x-1\]\[x=0:-A=-1\implies A=1\]\[x=\frac12:\frac12B=\frac32\implies B=3\]so we will your answer it looks like, which means we just have some simplification issue I guess

OpenStudy (anonymous):

Hmm, I still cant find the error. Your above answer is exactly how it went. Resulting in: 1/x + 3/(2x-1)

OpenStudy (anonymous):

Isnt int of (3/(2x-1)) --> 3*ln(2x-1) or am I stupid? :x

OpenStudy (turingtest):

you forgot the u-sub again, just like last time :)

OpenStudy (anonymous):

I still dont get that part :D

OpenStudy (turingtest):

\[\int{3\over 2x-1}dx\]\[u=2x-1\implies du=2dx\implies dx=\frac12du\]so we get\[\frac32\int{du\over u}\]

OpenStudy (anonymous):

Unless you're supposed to u-sub everytime you got a constant - not 1 - infront of the x.

OpenStudy (turingtest):

yes, you always must make sure you have the form du/u before integrating without the u-sub you don't have the right form

OpenStudy (anonymous):

Oh man! I shall be more careful in the future, apparently missed something major like this. Super thanks for your help TuringTest!

OpenStudy (turingtest):

Anytime :D